OpenStudy (cutiecomittee123):

Let logP/N=5 and logM/N=9 what is the relationship between P and M

OpenStudy (gibbs):

Subtract log M/ N by log P/N

OpenStudy (gibbs):

ok logMN−logPN=9−5 Can you simplify this a bit? maybe the right side

OpenStudy (cutiecomittee123):

4

OpenStudy (cutiecomittee123):

I get 4

OpenStudy (gibbs):

According to subtraction property of log logA−logB=logAB

OpenStudy (cutiecomittee123):

Okay

OpenStudy (cutiecomittee123):

here are the options I am given M=10000P M=1.0001P P=10000M M=4P

OpenStudy (cutiecomittee123):

So then the answer is M=4P??

OpenStudy (gibbs):

M = 10000p thats your answer

OpenStudy (cutiecomittee123):

how???

OpenStudy (gibbs):

log M/P = 4 log10MP=4 MP=104 ?

OpenStudy (gibbs):

elow?

OpenStudy (cutiecomittee123):

alright

OpenStudy (anonymous):

\[\log \frac{ M }{ N }-\log \frac{ P }{ N }=9-5\] \[\log \frac{ \frac{ M }{ N } }{ \frac{ P }{ N } }=4\] \[\frac{ M }{ P }=10^4=10000\] M=10000P